OpenStudy (anonymous):

The cost in dollars, C(d), to rent a houseboat during July to August from a certain company is given by C(d) = 120d + 100, where d is the number of rental days. a) Why does the 100 represent the fixed cost and the independent variable of the number of days? b) State the domain and range: Why is the domain: 0 ≤ d ≤ 62, range: 100 ≤ C ≤ 7540?

OpenStudy (anonymous):

a) independent means that it will not be changed no matter what, so in every case no matter how many days you rent it. you will always have to add 100 . that is why its independent.

OpenStudy (anonymous):

Because when I graphed the function it is no where 62 and 7540. But I do know that from July to Agusut = 62 days and 120(62)+100 = 7540.

OpenStudy (anonymous):

ohh, i see.

OpenStudy (anonymous):

july and aug have 31 days each so together its 62 days. so you can rent for a range between 0 adn 62 days.when you plug in those numbers you get the y values

OpenStudy (anonymous):

what do you mean when I plug in those numbers?

OpenStudy (anonymous):

well in a simple equation we have y=x+.... so the 0 to 62 are the numbers you would put in for the x and finish the equationto find the range of the y.

OpenStudy (anonymous):

i still don't get it :(

OpenStudy (anonymous):

in your equation:C(d) = 120d + 100, the d is the"x" and the C(d) is the y. so when you put in 0 you have C(d)=120*0+100. and you get C(d)=100. then plug in 62 day.: C(d)=120*62+100.cC(d)=7540::) so the smallest amoiunt was 100 and the largest amount was 7540

OpenStudy (anonymous):

ohhhh, i see. But how about x?

OpenStudy (anonymous):

so x represnts thedays you can rent. as i said before:how many days can someone rent?

OpenStudy (anonymous):

62 days so from 0 to 62?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ohh, thank you :)

OpenStudy (anonymous):

np